Solve
x -3/4 - x/7 = 5

Answer by stanbon(75887)

x -3/4 - x/7 = 5
-------
(6/7)x = 5 3/4
(6/7)x = 23/4
x = (7/6)(23/4)
x = 161/24
==================
Cheers,
Stan H.